OpenStudy (mama2dylan):

Basketball Suppose you make 90% of your free throws and you attempt 3 free throws. Use the Binomial Theorem to calculate each probability. a. You do not make any of them. b. You only make 1 of them. c. You only make 2 of them. d. You make all of them.

OpenStudy (amistre64):

well, what part of the thrm is losing you?

OpenStudy (amistre64):

(p+q)^3 = P(0) + P(1) + P(2) + P(3) p + q = 1, and they give you p=.90